Circular No. 8795

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 A3 (PETRIEW)

F. Fratev, E. Mihaylova, and A. Kirchev, Sofia, Bulgaria, report the recovery of comet P/2001 Q2 (cf. IAUC 7686) on unfiltered CCD images taken with a 0.25-m f/3.2 reflector at Zvezdno Obshtestvo Observatory (Plana), noting a diffuse coma of diameter ∼ 0′.2. Additional CCD astrometry has been reported by L. Buzzi and F. Luppi (Varese, Italy, 0.60-m reflector; comet ≈ 1′ east of prediction; quite diffuse with a round inner coma of size ∼ 15′′–20′′ in stacked 15-s exposures), W. Hasubick (Buchloe, Germany), and K. Kadota (Ageo, Saitama-ken, Japan, 0.25-m reflector; coma diameter 0′.7; diffuse with moderate condensation and no tail).

2007 UT α_{2000} δ_{2000} Mag. Observer
Jan. 11.69458 21^{h}15^{m}04.^{s}98 −9^{o}04′27′′.9 16.1 Fratev
11.69961 21 15 06.11 −9 04 20.7 16.1  _{”}
11.70273 21 15 06.61 −9 04 21.2 16.1  _{”}
13.70378 21 22 24.07 −8 43 33.7 15.9  _{”}
13.71309 21 22 26.41 −8 43 24.8 15.9  _{”}
13.71569 21 22 27.03 −8 43 24.3 15.9  _{”}
13.71957 21 22 27.88 −8 43 24.3  Buzzi
13.72247 21 22 28.45 −8 43 23.6  _{”}
13.72301 21 22 28.57 −8 43 16.1 15.9 Fratev
13.72495 21 22 29.04 −8 43 21.5  Buzzi
13.72621 21 22 29.33 −8 43 21.3 17.3 Hasubick
13.72713 21 22 29.33 −8 43 20.5 15.9 Fratev
13.72734 21 22 29.57 −8 43 19.4  Buzzi
13.73018 21 22 30.20 −8 43 17.6 16.5  _{”}
14.37541 21 24 52.63 −8 36 22.8  Kadota
14.37720 21 24 53.03 −8 36 21.5  _{”}
14.37906 21 24 53.46 −8 36 20.6 15.1  _{”}

The indicated correction to the prediction on MPC 51822 is ΔT = −0.04 day. The following orbital elements are by B. G. Marsden, from 434 observations 2001–2007 (mean residual 0′.7):

Epoch = 2007 Mar. 1.0 TT

T = 2007 Feb. 24.5806 TT  ω = 181.^{o}9215
e = 0.698129  Ω = 214.1025 }2000.0
q = 0.937602 AU  i = 13.9747

a = 3.105970 AU n^{o} = 0.1800563 P = 5.47 years

2007 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green